OpenStudy (ilovebmth1234):

What is the mass of 0.8 moles of Iron? A) 12.1 grams B) 25.2 grams C) 44.7 grams D) 100.8 grams

OpenStudy (ramirocruzo):

Mass=Number of moles X Atomic Mass or Molecular Mass Atomic Mass of Fe= 56u

OpenStudy (ilovebmth1234):

i saw that on the yahoo answers but the answer has to be A,B,C,or D

OpenStudy (ramirocruzo):

???? I didnt copy...... (T~T) Its C) 44.7gm

OpenStudy (ilovebmth1234):

no im sorry i didnt think you copied lol i was just saying i saw that on a yahoo question :) im sorry and thanks would you like to help me with one more?

OpenStudy (ramirocruzo):

Yup....

OpenStudy (ramirocruzo):

???

OpenStudy (ilovebmth1234):

Using the Periodic Table, determine the atomic mass of Na2C2O4. A) 134 grams B) 124 grams C) 114 grams D) 104 grams

OpenStudy (ramirocruzo):

Just solve it yourself... :p 2Na+2C+4O=2(23)+2(12)+4(16)=??? Note: Its molecular mass not atomic mass

OpenStudy (whpalmer4):

Add up the atomic mass of two sodium atoms, two carbon atoms, and four oxygen atoms. That's your answer. Should be able to do these on your own with the knowledge you have now.

OpenStudy (ramirocruzo):

Hmmmmm,,, It comes up to be 134gms....

OpenStudy (ilovebmth1234):

thanks guys and @whpalmer4 thanks for breaking it down i appreciate both of you